OpenStudy (anonymous):

CAN ANYONE PLZZZ HELP ME!!!!!!!!!! find the value of 10√2/√5

OpenStudy (anonymous):

10√2/√5 =>5*2sqrt2/sqrt5 =>can be written as sqrt5*sqrt5*2*sqrt2/sqrt5 =>2sqrt10 (Ans)

OpenStudy (anonymous):

omg thank you!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

\[10\sqrt{2}/\sqrt{5}=\sqrt{(10\sqrt{2}/\sqrt{5})^2}=\sqrt{10^2*\sqrt{2}^2/\sqrt{5}^2}=\sqrt{100*2/5}=\sqrt{40}\]\[=2\sqrt{10}\approx 6.32455532034\]

OpenStudy (anonymous):

can you help me out with one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

explain how youy find the exact value of 3/√8

OpenStudy (anonymous):

You can write 3/sqrt8 as 3/√(2*2*2) =>3/2√2 =>3/2 = 1.5 You can further divide 1.5 by √2 √2=1.414 (You have to remember this Then do the division 1.5/1.414

OpenStudy (anonymous):

ok so the answer would be 1.06